# Mechanistic Modeling of Continuous Lyophilization for Pharmaceutical Manufacturing

### v1

@@ Conclusion
This work proposes the first mechanistic model for the suspended-vial continuous lyophilization technology, in which vials containing the pharmaceutical products are suspended and moved continuously along the process. The model can be used to simulate the entire lyophilization process, including the freezing, primary drying, and secondary drying steps. The freezing model considers preconditioning, vacuum-induced surface freezing (VISF), spontaneous and controlled nucleation, solidification, and final cooling. The primary drying model captures heat transfer in the frozen region and ice sublimation driven by mass transfer at the sublimation front. The secondary drying model describes simultaneous heat transfer in the dried region and bound water desorption. The overall model can describe the evolution of product temperature, ice/water fraction, and residual moisture throughout the process.
Our model is validated using experimental data from the literature, in which the model predictions are consistent with the experimental measurements in all cases. The validated model is applied for four case studies that entail continuous lyophilization process design and optimization, detailed investigation of VISF, model modifications for conventional drying, and condenser failure analysis. The framework and results presented in this work are promising and could be used to guide the design and development of future continuous lyophilization technology.
With a well-developed mechanistic model, future work could consider utilizing the model for several crucial tasks, e.g., uncertainty quantification, state estimation, and control design, which will ultimately lead to the development of a high-quality digital twin for continuous lyophilization that can greatly facilitate the process design and development in this area.
@@
All software and data will be made available upon publication of the manuscript.
Acknowledgements
This research was supported by the U.S. Food and Drug Administration under the FDA BAA-22-00123 program, Award Number 75F40122C00200.
@@ Acknowledgements
Appendix A Supplemental equations
@@
This appendix summarizes additional equations that are not crucial part of the model but required for calculating some model parameters.
@@
$\rho_{\textrm{l}}=\frac{V_{\textrm{l}}}{\dfrac{x_{\textrm{s}}}{\rho_{\textrm{s}}}+\dfrac{1-x_{\textrm{s}}}{\rho_{\textrm{w}}}}.$ (A.1)
@@
$m_{\textrm{s}}=x_{\textrm{s}}\rho_{\textrm{l}}V_{\textrm{l}},$ (A.2)
$m_{\textrm{w,0}}=(1-x_{\textrm{s}})\rho_{\textrm{l}}V_{\textrm{l}}.$ (A.3)
Then, (1) can be used as usual.
During the freezing step, the surface area $A_{r}$ changes with time due to changes in the total volume caused by phase transition. In general, the surface area can be calculated by
$A_{r}(t)=\frac{4\!\left(\dfrac{m_{\textrm{s}}}{\rho_{\textrm{s}}}+\dfrac{m_{\textrm{w}}(t)}{\rho_{\textrm{w}}}+\dfrac{m_{\textrm{i}}(t)}{\rho_{\textrm{i}}}\right)}{d},$ (A.4)
@@
$V_{\textrm{f}}=\pi d^{2}(H-S).$ (A.5)
The properties of the frozen region (subscript ‘f’) can be calculated from those of the ice and solid material or solute as follows:
$\rho_{\textrm{f}}=\frac{1}{\dfrac{x_{\textrm{s}}}{\rho_{\textrm{s}}}+\dfrac{1-x_{\textrm{s}}}{\rho_{\textrm{i}}}},$ (A.6)
$C_{p,\textrm{f}}=x_{\textrm{s}}C_{p,\textrm{s}}+(1-x_{\textrm{s}})C_{p,\textrm{i}},$ (A.7)
$k_{\textrm{f}}=x_{\textrm{s}}k_{\textrm{s}}+(1-x_{\textrm{s}})k_{\textrm{i}}.$ (A.8)
Here we assume that the final mass of ice is equal to the initial mass of water (complete freezing), meaning that the amount of water loss due to evaporation during VISF and unfrozen water (bound water) is negligible, which is generally true.
Finally, the height/thickness of the product $H$ can be calculated by
$H=\frac{(m_{\textrm{s}}+m_{\textrm{w,0}})}{\rho_{\textrm{f}}A_{z}}.$ (A.9)
Note that, for some experiments that focus on primary drying, $\rho_{\textrm{f}},C_{p,\textrm{f}},k_{\textrm{f}},$ and $H$ are reported, and so the above calculations may not be needed.
@@
$V_{\textrm{e}}=\pi d^{2}H,$ (A.10)
which is constant because there is no moving boundary in this case.
Appendix B Numerical methods
This appendix comprehensively describes numerical techniques required for solving the model equations and summarizes the final equations implemented in MATLAB.
For the freezing model, since a lumped thermal capacity approximation is used, the resulting equations are all ODEs, and so the equations can be solved directly with the selected ODE solver, which is ode15s in this case.
For the primary drying model, the PDE (28) is discretized spatially to produce a system of ODEs, which is then fed to the ODE solver; the technique is referred to as the method of lines [41]. Since the PDE (28) involves a moving interface $S(t)$, define a new variable
$\xi=\frac{z-S}{H-S},$ (B.1)
where $\xi$ represents is dimensionless position with respect to the moving interface (aka sublimation front). This dimensionless position varies from 0 to 1, in which $\xi=0$ at $z=S$ and $\xi=1$ at $z=H$. Consequently, the PDE (28) and boundary conditions (34) and (35) become
$\rho_{\textrm{f}}C_{p,\textrm{f}}\dfrac{\partial T}{\partial t}=\frac{k_{\textrm{f}}}{(H-S)^{2}}\dfrac{\partial^{2}T}{\partial\xi^{2}}-\frac{\rho_{\textrm{f}}C_{p,\textrm{f}}(\xi-1)}{H-S}\frac{dS}{dt}\frac{\partial T}{\partial\xi}+\frac{Q_{\textrm{rad}}}{V_{\textrm{f}}},\qquad 0<\xi<1,$ (B.2)
$-\frac{k_{\textrm{f}}}{H-S}\dfrac{\partial T}{\partial\xi}=h_{\textrm{b}}(T-T_{\textrm{b}}),\qquad\xi=1,$ (B.3)
$N_{\textrm{w}}\Delta H_{\textrm{sub}}=\frac{k_{\textrm{f}}}{H-S}\dfrac{\partial T}{\partial\xi}+\sigma\varepsilon_{\textrm{gl}}\left(T_{\textrm{c}}^{4}-T^{4}\right),\qquad\xi=0.$ (B.4)
@@ Appendix B Numerical methods
$\Delta\xi=\frac{1}{n_{z}-1},$ (B.5)
$\xi=j\Delta\xi,$ (B.6)
where $n_{z}$ is the number of grid points, $\Delta\xi$ is the distance between each grid point, and $j$ is the integer index for discretization, in which $j=1$ at $z=S$ and $j=n_{z}$ at $z=H$. Discretizing (B.2)–(B.4) using the second-order finite difference scheme results in
$\begin{split}\rho_{\textrm{f}}C_{p,\textrm{f}}\dfrac{dT_{j}}{dt}=\frac{k_{\textrm{f}}}{(H-S)^{2}}\!\left(\dfrac{T_{j+1}-2T_{j}+T_{j-1}}{\Delta\xi^{2}}\right)\!-\frac{\rho_{\textrm{f}}C_{p,\textrm{f}}(j\Delta\xi-1)}{H-S}\frac{dS}{dt}\!\left(\dfrac{T_{j+1}-T_{j-1}}{2\Delta\xi}\right)\!+\frac{Q_{\textrm{rad}}}{V_{\textrm{f}}},\\ \qquad j=1,2,\dots,n_{z},\end{split}$ (B.7)
$-\frac{k_{\textrm{f}}}{H-S}\!\left(\dfrac{T_{j+1}-T_{j-1}}{2\Delta\xi}\right)\!=h_{\textrm{b}}(T_{j}-T_{\textrm{b}}),\qquad j=n_{z},$ (B.8)
$N_{\textrm{w}}\Delta H_{\textrm{sub}}=\frac{k_{\textrm{f}}}{H-S}\!\left(\dfrac{T_{j+1}-T_{j-1}}{2\Delta\xi}\right)\!+\sigma\varepsilon_{\textrm{gl}}\!\left(T_{\textrm{c}}^{4}-T_{j}^{4}\right),\qquad j=1,$ (B.9)
where $T_{j}$ is the product temperature at position $j$. The ghost point at $j=n_{z}+1$ in (B.8) can be eliminated by substituting $T_{j=n_{z}+1}$ into the governing equation (B.7) for $j=n_{z}$. Similarly, $T_{j=0}$ in (B.9) can be eliminated by substituting $T_{j=n_{z}+1}$ into the governing equation (B.7) for $j=1$. The above nondimensionalization and discretization techniques result in a system of nonlinear ODEs defined on a moving-grid system, which is then integrated using ode15s.
Numerical treatment for the secondary drying model are simpler than those described for the primary drying model as there is no moving interface. Define
$\Delta z=\frac{H}{n_{z}-1},$ (B.10)
$z=j\Delta z,$ (B.11)
where $\Delta z$ is the distance between each grid point. Discretizing (38), (39), (41), and (42) using the second-order finite difference scheme results in
$\frac{dT_{j}}{dt}=\frac{k_{\textrm{e}}}{\rho_{\textrm{e}}C_{p,\textrm{e}}}\!\left(\frac{T_{j+1}-2T_{j}+T_{j-1}}{\Delta z^{2}}\right)\!+\frac{\rho_{d}\Delta H_{\textrm{des}}}{\rho_{\textrm{e}}C_{p}}\!\left(\dfrac{dc_{\textrm{w},j}}{dt}\right)\!+\frac{Q_{\textrm{rad}}}{V_{\textrm{e}}},\qquad j=1,2,\dots,n_{z},$ (B.12)
$\dfrac{dc_{\textrm{w},j}}{dt}=f_{\textrm{a}}e^{-E_{\textrm{a}}/RT_{j}}c_{\textrm{w},j},\qquad j=1,2,\dots,n_{z},$ (B.13)
$-k_{\textrm{e}}\!\left(\dfrac{T_{j+1}-T_{j-1}}{2\Delta z}\right)\!=h_{\textrm{b}}(T_{j}-T_{\textrm{b}}),\qquad j=n_{z},$ (B.14)
$-k_{\textrm{e}}\!\left(\dfrac{T_{j+1}-T_{j-1}}{2\Delta z}\right)\!=\sigma\varepsilon_{\textrm{gl}}\!\left(T_{\textrm{c}}^{4}-T_{j}^{4}\right),\qquad j=1,$ (B.15)
@@
$j_{\textrm{w,max}}$ maximum mass flow rate of a condenser (kg/s)
@@
$m_{\textrm{i,n}}$ mass of ice resulting from the first nucleation (kg)
@@
$T_{\textrm{b}}$ bottom shelf temperature (K)
$T_{\textrm{c}}$ chamber wall temperature (K)
$T_{\textrm{f}}$ freezing point (K)
$T_{\textrm{g}}$ cryogenic gas temperature (K)
$T_{\textrm{n}}$ nucleation temperature (K)
@@
$t_{\textrm{0}}$ initial time, (s)
@@
c chamber (environment)

### v2

@@ Conclusion
This article proposes the first mechanistic model for continuous lyophilization, in which the state-of-the-art suspended-vial technology is considered. The model can simulate the entire lyophilization process, including the freezing, primary drying, and secondary drying steps. The freezing model considers preconditioning, vacuum-induced surface freezing (VISF), spontaneous and controlled nucleation, solidification, and cooling. The primary drying model captures heat transfer in the frozen region and mass transfer resulting from sublimation of ice crystals at the sublimation front. The secondary drying model describes simultaneous heat transfer in the dried region and desorption of bound water. The overall model can describe the evolution of product temperature, ice/water fraction, and residual moisture throughout the process.
The proposed model is validated for all three steps including freezing, primary drying, and secondary drying, in which the model predictions are consistent with the experimental measurements in all cases. The validated model is demonstrated for a variety of applications related to process design and optimization. Every single simulation can be run within less than 1 s on a normal laptop, allowing the model to be employed for any purpose. The framework and results presented in this work are suitable for guiding the design and development of future continuous lyophilization technology.
With a well-developed mechanistic model, future work could consider utilizing the model for several important tasks, e.g., state estimation, optimal control design, and uncertainty quantification, which will ultimately support the development of a high-quality digital twin for continuous lyophilization to advance the manufacturing process.
@@
All software and data will be made available upon publication of the manuscript. Both MATLAB and Julia implementations will be provided.
Acknowledgments
RDB and PS were supported by the U.S. Food and Drug Administration under the FDA BAA-22-00123 program, Award Number 75F40122C00200. The authors would like to thank Pitipat Wongsittikan for helping optimize part of the Julia code.
@@ Acknowledgments
Appendix A Analysis of transport phenomena
Biot number and spatial dimensions
As discussed in the main text, the freezing step is modeled using the lumped capacity method, whereas the drying steps are analyzed in one spatial dimension, a vertical direction. Various studies have shown via simulation that variation in the radial direction is negligible [17, 13]. This section provides a detailed analysis from the theoretical perspective to justify this assumption and briefly discusses strategies to modify the model for situations where this assumption is not true.
Introduce the Biot number
$\textrm{Bi}=\frac{hL}{k},$ (A.1)
where $L$ is the characteristic length or length scale, $h$ is the heat transfer coefficient, and $k$ is the thermal conductivity. In general, the length scale is defined as the ratio of the volume to the surface area, which results in $r_{\textrm{o}}/2$ for a long cylinder. However, for conservative analysis, it is recommended to use the actual radius $r_{\textrm{o}}$ instead [45]. The Biot number is defined as the ratio of the internal conduction resistance to the external convection resistance, and so a small Bi value implies that the resulting temperature gradient should also be small [31]. In various heat transfer textbooks, the value of 0.1 is recommended as a guideline; if Bi $<0.1$, a lumped capacity model can be used. However, in practice, a lumped capacity model could still be valid even for Bi $>0.1$, depending on the system, which is analyzed here.
[…]
Figure A.1: Comparison between the 1D and lumped capacity models for the (A) volume-averaged temperature and (B) total heat removed for the default model parameters.
[…]
Figure A.2: Comparison between the 1D and lumped capacity model for the (A) volume-averaged temperature and (B) total heat removed for the conservative case.
Finally, the errors (differences) between both models for different values of Bi are shown in Fig. A.3. Of course, if the Biot is too high, the lumped capacity model will not be accurate. That situation could happen, e.g., when a very large vial is used or the heat transfer coefficient is much higher than usual. In such cases, a high-fidelity model might be needed [13]. However, the analysis in this section suggests that a lumped capacity model is highly accurate for typical lyophilization applications, in particular for unit doses.
Figure A.3: Errors between the 1D and lumped capacity model for the (A) volume-averaged temperature and (B) total heat removed for different Bi.
Mass transport in the dried product
In secondary drying, the two important mechanisms for mass transport are (1) vapor phase or water vapor transport through the porous dried product and (2) desorption of bound water from the solid surface of the dried product. Previous simulation and experimental data suggest that desorption is the rate-limiting step [37, 23, 4, 24, 25, 36], and so the equations/terms related to vapor phase transport are typically omitted from recently published models, including our model. This section provides an analysis from the perspective of transport phenomena to justify this assumption.
First, consider the vapor phase transport through the porous dried product. Modeling the transport of liquids and gases in porous media is challenging as it could entail a variety of mechanisms besides ordinary diffusion (e.g., Knudsen diffusion), and thus no fully satisfactory theory currently exists [46]. Nevertheless, relative success has been achieved by using the concept of an effective diffusivity; in this analysis, we follow the procedure described in Ref. [31]. Define the effective diffusivity $D_{\textrm{e}}$ by
$\frac{1}{D_{\textrm{e}}}=\frac{1}{D_{\textrm{e,g}}}+\frac{1}{D_{\textrm{e,K}}},$ (A.2)
where $D_{\textrm{e,g}}$ is the effective diffusion coefficient for ordinary diffusion, i.e., diffusion of the molecules in the vapor/gas phase, $D_{\textrm{e,K}}$ is the effective diffusion coefficient for Knudsen diffusion, and $D_{\textrm{e}}$ is the effective diffusivity that combines the effects of ordinary diffusion and Knudsen diffusion. For ordinary diffusion,
$D_{\textrm{e,g}}=\frac{\epsilon}{\tau}D_{\textrm{g}},$ (A.3)
where $D_{\textrm{g}}$ is the binary diffusion coefficient, e.g., water vapor in air, $\epsilon$ is the porosity, and $\tau$ is the tortuosity. For Knudsen diffusion, a similar expression can be used, that is,
$D_{\textrm{e,K}}=\frac{\epsilon}{\tau}D_{\textrm{K}},$ (A.4)
where $D_{\textrm{K}}$ is the Knudsen diffusion coefficient. This coefficient could be calculated from
$D_{\textrm{K}}=97\bar{r}\sqrt{\frac{T}{M}},$ (A.5)
where $\bar{r}$ is the average pore radius, $T$ is the temperature, and $M$ is the molar mass in g/mol.
With the effective diffusivity calculated by (A.2)–(A.5), the total mass flux can be calculated using Fick’s law of diffusion, with the binary diffusion coefficient replaced by the effective diffusivity. Then, a continuity (mass balance) equation for the water vapor can be written and coupled to the mechanistic model proposed in the main paper. The key benefit of this technique is that the implementation closely resembles the approach for solving problems entailing ordinary diffusion of a binary mixture, which is relatively straightforward. Nevertheless, additional information is needed, including the porosity, tortuosity, and average pore radius. A more complicated strategy that is used in the literature is to apply the dusty gas model, which is derived from the Maxwell-Stefan Equation (see examples in [16, 19]). This approach, however, requires many more parameters and equations.
Next, we compare the significance of vapor phase transport and desorption. Scale analysis can be used to analyze the contribution of each transport process (see various examples in [31]). Properly estimating the time scale of each phenomenon can help identify the key process that should be considered in the model and omit unimportant phenomena. The time scale (aka time constant) for mass diffusion can be approximated by [47, 31]
$t_{\text{dff}}=\frac{L^{2}}{D_{\textrm{e}}},$ (A.6)
where $t_{\text{dff}}$ is the time scale for diffusion and $L$ is the proper length scale. Similarly, the time scale for desorption can be approximated by
$t_{\text{des}}=\frac{1}{k_{\textrm{d}}},$ (A.7)
where $t_{\text{des}}$ is the time scale for desorption and $k_{\textrm{d}}$ is the desorption constant defined in (43). By comparing $t_{\text{dff}}$ and $t_{\text{des}}$, the limiting step for mass transfer in the secondary drying step can be identified, without any simulation or experimental data.
For numerical comparison, a set of parameter values are obtained from Refs. [15, 17], namely $\bar{r}=5$$\times$$10^{-6}$ m, $\epsilon=0.815$, $\tau=1.2$, and $k_{\textrm{d}}=7.8$$\times$$10^{-5}$ s-1. The binary coefficient of water in air is obtained from Ref. [31], which is $D_{\textrm{g}}=1.97$$\times$$10^{-5}$ m2/s at 256 K and 1 atm. In secondary drying, the operating pressure is much lower than 1 atm, and so the actual $D_{\textrm{g}}$ is lower as the diffusion coefficient of gas mixtures at low pressure is inversely proportional to the pressure [31, 46], which makes our analysis more conservative. The length scale for this analysis is set to 0.01 m, a typical vial/sample size (thickness/diameter). Substituting these values into (A.6) and (A.7) gives $t_{\text{dff}}=7.5$ s and $t_{\text{des}}=3.6$ h. Obviously, the time scale for diffusion is much smaller than for desorption by several orders of magnitude, indicating that desorption is the rate-limiting step. A typical drying time of secondary drying is on the order of several hours, agreeing with the calculated $t_{\text{des}}$ of about 3.6 h. On the other hand, the time scale for diffusion is only 7.5 s, meaning that diffusion is much faster and so can be safely omitted from the model. Even some of the parameters are changed by an order of magnitude, this conclusion does not change.***This analysis holds for lyophilization of unit doses only; for lyophilization of bulk material (e.g., in food applications), the length scale of the product could be much larger, and so the conclusion might be different.
Finally, there is another mass transfer mechanism that is not commonly discussed or modeled, namely diffusion of water in the solid matrix (not in the pores), i.e., solid-phase diffusion [37]. Diffusion in the solid matrix is very slow, and so it could limit the mass transport in secondary drying. Modeling this phenomenon is challenging as it requires detailed knowledge of the solid matrix, particle size, and locations of bound water in the matrix, which is generally not available. A similar approximation to (A.6) can be carried out. By using the reported length scale of about $5$$\times$$10^{-7}$ m and diffusion coefficient of water in the solid phase of about $7$$\times$$10^{-16}$ m2/s [37], the time scale for solid-phase diffusion is about 10 min, which is still much lower than for desorption.
To summarize, this section details the scale analysis that can be efficiently used to evaluate the effects of various mass transfer mechanisms. For lyophilization of pharmaceuticals in unit doses, our analysis shows that desorption of bound water is the rate-limiting step for mass transfer, and so other mechanisms can be omitted from the model, agreeing with both simulation and experimental data in the literature.
Appendix B Supplemental equations
@@
This appendix summarizes equations that are not a crucial part of the model but required for calculating some model parameters.
@@
$\rho_{\textrm{l}}=\frac{V_{\textrm{l}}}{\dfrac{x_{\textrm{s}}}{\rho_{\textrm{s}}}+\dfrac{1-x_{\textrm{s}}}{\rho_{\textrm{w}}}}.$ (B.1)
@@
$m_{\textrm{s}}=x_{\textrm{s}}\rho_{\textrm{l}}V_{\textrm{l}},$ (B.2)
$m_{\textrm{w,0}}=(1-x_{\textrm{s}})\rho_{\textrm{l}}V_{\textrm{l}}.$ (B.3)
Then (1) can be used as usual.
During the freezing step, the surface area $A_{r}$ varies with time due to changes in the total volume caused by phase transition. In general, the surface area can be calculated by
$A_{r}(t)=\frac{4\!\left(\dfrac{m_{\textrm{s}}}{\rho_{\textrm{s}}}+\dfrac{m_{\textrm{w}}(t)}{\rho_{\textrm{w}}}+\dfrac{m_{\textrm{i}}(t)}{\rho_{\textrm{i}}}\right)}{d},$ (B.4)
@@
$V_{\textrm{f}}=\pi d^{2}(H-S).$ (B.5)
The height/thickness of the product $H$ can be calculated by
$H=\frac{m_{\textrm{s}}+m_{\textrm{w,0}}}{\rho_{\textrm{f}}A_{z}}.$ (B.6)
The properties of the frozen region (subscript ‘f’) can be calculated from those of the ice and solid material or solute from
$\rho_{\textrm{f}}=\frac{1}{\dfrac{x_{\textrm{s}}}{\rho_{\textrm{s}}}+\dfrac{1-x_{\textrm{s}}}{\rho_{\textrm{i}}}},$ (B.7)
$C_{p,\textrm{f}}=x_{\textrm{s}}C_{p,\textrm{s}}+(1-x_{\textrm{s}})C_{p,\textrm{i}},$ (B.8)
$k_{\textrm{f}}=x_{\textrm{s}}k_{\textrm{s}}+(1-x_{\textrm{s}})k_{\textrm{i}}.$ (B.9)
These expressions assume that the final mass of ice is equal to the initial mass of water (complete freezing), meaning that the amount of water loss due to evaporation during VISF and unfrozen water (bound water) is negligible, which is generally true. Note that some experimental primary drying studies report $\rho_{\textrm{f}},C_{p,\textrm{f}},k_{\textrm{f}},$ and/or $H$, in which case the above calculations may not be needed.
@@
$V_{\textrm{e}}=\pi d^{2}H,$ (B.10)
which is constant because there is no moving boundary (i.e., sublimation front) in this case.
Appendix C Numerical methods
This appendix comprehensively describes numerical techniques required for solving the model equations efficiently. With the given procedure, our model can be simulated accurately within 1 s on a normal laptop.
For the freezing model, since the lumped capacity approximation is used, the resulting equations are ordinary differential equations (ODEs), and so the equations can be integrated directly with a proper ODE solver. This work uses MATLAB’s ode15s.
For PDEs in the primary drying and secondary drying models, the method of lines [48] is recommended. The method of lines consists of two steps. First, the PDEs are discretized spatially to produce a system of ODEs. Second, the resulting ODEs are integrated with a proper ODE solver (MATLAB’s ode15s in this work).
For the primary drying model, first consider the PDE (32). As the equation involves a moving interface $S(t)$, define a new variable
$\xi=\frac{z-S}{H-S},$ (C.1)
where $\xi$ is the dimensionless position with respect to the moving interface (aka sublimation front). This dimensionless position varies from 0 to 1, in which $\xi=0$ at $z=S$ and $\xi=1$ at $z=H$. Consequently, the PDE (32) and boundary conditions (38) and (39) become
$\rho_{\textrm{f}}\,C_{p,\textrm{f}}\,\dfrac{\partial T}{\partial t}=\frac{k_{\textrm{f}}}{(H-S)^{2}}\dfrac{\partial^{2}T}{\partial\xi^{2}}-\frac{\rho_{\textrm{f}}\,C_{p,\textrm{f}}\,(\xi-1)}{H-S}\frac{dS}{dt}\frac{\partial T}{\partial\xi}+\frac{Q_{\textrm{rad}}}{V_{\textrm{f}}},\qquad 0<\xi<1,$ (C.2)
$-\frac{k_{\textrm{f}}}{H-S}\dfrac{\partial T}{\partial\xi}=h_{\textrm{b}}(T-T_{\textrm{b}}),\qquad\xi=1,$ (C.3)
$N_{\textrm{w}}\Delta H_{\textrm{sub}}=\frac{k_{\textrm{f}}}{H-S}\dfrac{\partial T}{\partial\xi}+\sigma\mathcal{F}_{\textrm{s1}}(T_{\textrm{u}}^{4}-T^{4}),\qquad\xi=0.$ (C.4)
@@ Appendix C Numerical methods
$\Delta\xi=\frac{1}{n_{z}-1},$ (C.5)
$\xi=(j-1)\Delta\xi,$ (C.6)
where $n_{z}$ is the number of grid points, $\Delta\xi$ is the distance between each grid point, and $j$ is the integer index for discretization, in which $j=1$ at $z=S$ and $j=n_{z}$ at $z=H$. Discretizing (C.2)–(C.4) using the second-order finite difference scheme results in
$\begin{split}\dfrac{dT_{j}}{dt}=\frac{k_{\textrm{f}}}{\rho_{\textrm{f}}C_{p,\textrm{f}}(H-S)^{2}}\!\left(\dfrac{T_{j+1}-2T_{j}+T_{j-1}}{\Delta\xi^{2}}\right)\!-\frac{j\Delta\xi-1}{H-S}\frac{dS}{dt}\!\left(\dfrac{T_{j+1}-T_{j-1}}{2\Delta\xi}\right)\!+\frac{Q_{\textrm{rad}}}{\rho_{\textrm{f}}C_{p,\textrm{f}}V_{\textrm{f}}},\\ \qquad j=1,2,\dots,n_{z},\end{split}$ (C.7)
$-\frac{k_{\textrm{f}}}{H-S}\!\left(\dfrac{T_{j+1}-T_{j-1}}{2\Delta\xi}\right)\!=h_{\textrm{b}}(T_{j}-T_{\textrm{b}}),\qquad j=n_{z},$ (C.8)
$N_{\textrm{w}}\Delta H_{\textrm{sub}}=\frac{k_{\textrm{f}}}{H-S}\!\left(\dfrac{T_{j+1}-T_{j-1}}{2\Delta\xi}\right)\!+\sigma\mathcal{F}_{\textrm{s1}}\!\left(T_{\textrm{u}}^{4}-T_{j}^{4}\right),\qquad j=1,$ (C.9)
where $T_{j}$ is the product temperature at position $j$. The ghost point at $j=n_{z}+1$ in (C.8) can be eliminated by substituting $T_{j=n_{z}+1}$ into the governing equation (C.7) for $j=n_{z}$. Similarly, $T_{j=0}$ in (C.9) can be eliminated by substituting $T_{j=0}$ into the governing equation (C.7) for $j=1$. The above nondimensionalization and discretization techniques result in a system of nonlinear ODEs defined on a moving-grid system.
Numerical treatment for the secondary drying model is simpler than those described for the primary drying model as there is no moving interface. Define
$\Delta z=\frac{H}{n_{z}-1},$ (C.10)
$z=(j-1)\Delta z,$ (C.11)
where $\Delta z$ is the distance between each grid point, $j=1$ at $z=0$, and $j=n_{z}$ at $z=H$. Discretizing (42), (43), (45), and (46) using the second-order finite difference scheme results in
$\frac{dT_{j}}{dt}=\frac{k_{\textrm{e}}}{\rho_{\textrm{e}}C_{p,\textrm{e}}}\!\left(\frac{T_{j+1}-2T_{j}+T_{j-1}}{\Delta z^{2}}\right)\!+\frac{\rho_{d}\Delta H_{\textrm{des}}}{\rho_{\textrm{e}}C_{p,\textrm{e}}}\!\left(\dfrac{dc_{\textrm{w},j}}{dt}\right)\!+\frac{Q_{\textrm{rad}}}{\rho_{\textrm{e}}C_{p,\textrm{e}}V_{\textrm{e}}},\qquad j=1,2,\dots,n_{z},$ (C.12)
$\dfrac{dc_{\textrm{w},j}}{dt}=-f_{\textrm{a}}e^{-E_{\textrm{a}}/RT_{j}}c_{\textrm{w},j},\qquad j=1,2,\dots,n_{z},$ (C.13)
$-k_{\textrm{e}}\!\left(\dfrac{T_{j+1}-T_{j-1}}{2\Delta z}\right)\!=h_{\textrm{b}}(T_{j}-T_{\textrm{b}}),\qquad j=n_{z},$ (C.14)
$-k_{\textrm{e}}\!\left(\dfrac{T_{j+1}-T_{j-1}}{2\Delta z}\right)\!=\sigma\mathcal{F}_{\textrm{s1}}\!\left(T_{\textrm{u}}^{4}-T_{j}^{4}\right),\qquad j=1,$ (C.15)
@@
$D$ diffusivity (m2/s)
$D_{\textrm{e}}$ effective diffusivity (m2/s)
$D_{\textrm{e,g}}$ effective diffusivity for ordinary diffusion of gases (m2/s)
$D_{\textrm{e,K}}$ effective diffusivity for Knudsen diffusion (m2/s)
$D_{\textrm{g}}$ binary diffusion coefficient of gases (m2/s)
$D_{\textrm{K}}$ Knudsen diffusion coefficient (m2/s)
@@
$j_{\textrm{w,max}}$ condenser maximum capacity (kg/s)
@@
$r_{\textrm{o}}$ radius of the vial (m)
$\bar{r}$ average pore radius (m)
@@
$t_{\textrm{0}}$ initial time (s)
@@
$t_{\textrm{dff}}$ time scale for diffusion (s)
$t_{\textrm{des}}$ time scale for desorption (s)
@@
c chamber (environment), chamber wall (side)